OpenStudy (anonymous):

solve by factoring. 5x^2+13x-46=0 two solutions:

OpenStudy (cwrw238):

factor 5 = 5*1 -46 = -2*23 or 2 * -23

OpenStudy (anonymous):

(5x + 23)(x - 2)=0 solutions x=2, x=-23/5

OpenStudy (anonymous):

(5x+23)(x-2) then x=2 or -23/5

OpenStudy (cwrw238):

from the factors u need to 2 numbers whose product is -46 and sum is + 13

OpenStudy (anonymous):

thanks!\

OpenStudy (cwrw238):

- sorry - you have to involve the 5 as well 5 * -2 + 23 = + 13